Mandy is now twice Katy's age. In six years she will be 1.5 times Katy's age.
How old is Katy

Answer by Paul(988)

Since Mandy is twice as old as cathy, you know that there is one variable to consider, because due supportment.
Kathy's age = x
Mandy's age = 2x
In six years Kathy's age = x+6
In six years Mandy's age = 2x+6
And since she will be 1.5 the quanity of Kathy in SIX years we write.
1.5(x+6) Expression 1
And Mandy herself is 2x+6 Expression 2.
Expression 1 = Expression 2.
1.5(x+6)=2x+6
1.5x+9=2x+6
0.5x=3
x=6 -----> Kathy's age.
2(6)=12 --> Mandy's age.

Check: Kathy's age in 6 years = 6+6=12
Mandy's age in 6 years = 12+6=18
And 12*1.5=18
or 18/1.5=12

Hence, Kathy is 6 years old, and Mandy is 12 years old.
Paul.